OpenStudy (anonymous):

solve the equation by completing the square. Round to the nearest hundredth if necessary. 3. x2 + 2x – 8 = 0 (1 point) 4, 2 –4, 2 4, –2 –4, –2

OpenStudy (hoblos):

x2 + 2x – 8 = 0 x2 +2x +1 =9 (x+1)^2 =9 x+1 = ±3 so x+1=3 or x+1=-3 x=2 x=-4